OpenStudy (anonymous):

Help Please ;) Which ratio forms a proportion 18 over 38

OpenStudy (anonymous):

Didnt you already post this?

OpenStudy (anonymous):

I got no help tho well i got help but it wasnt clear enough

OpenStudy (anonymous):

and the person left so i couldn't ask a question

OpenStudy (mathstudent55):

A proportion is a ratio equaling another ratio. 18/38 is one ratio. If you reduce 18/38 to 9/19, you can write 18/38 = 9/19. That's a proportion.

OpenStudy (anonymous):

ik but nun of the answer choses match that

OpenStudy (anonymous):

The one answer was 9/29, right?

OpenStudy (mathstudent55):

Is this all the question states? Where are the choices for answers?

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

\[\frac{ 7 }{ 27 }\frac{ 24 }{ 44 } \frac{ 9 }{ 29 } \frac{ 27 }{ 57 }\]

OpenStudy (anonymous):

Well i tried subtracting 9 from the top and bottom,and thats what i got..

OpenStudy (mathstudent55):

You have 18/38 = (9 * 2)/(19 * 2) = 9/19 The choices are: 7/27 = 7/27 (can't be reduced bec 7 is prime) 24/44 = (6 * 4)/(11 * 4) = 6/11 9/29 = 9/29 (can't be reduced bec 29 is prime) 27/57 = (9 * 3)/(19 * 3) = 9/19 Now compare the reduced 18/38 with each of the reduced choices.